Alan is a single taxpayer with a gross income of $88,000, a taxable income of $66,000,...

  1. Alan is a single taxpayer with a gross income of $88,000, a taxable income of $66,000, and an income tax liability of $10,460 for 2018. Alan also has $8,000 of tax-exempt interest income. What are Alan's marginal, average, and effective tax rates?
    1. 22% marginal; 14.5% average; 15.8% effective.
    2. 24% marginal; 14.5% average; 15.8% effective.
    3. 12% marginal; 15.8% average; 14.1% effective.
    4. 22% marginal; 15.8% average; 14.1% effective.
Answer #1

2019

taxable income
0-19400 10%
19401-78950 12%

As the taxable income 66000$ falls between 19401-78950 the marginal tax rates is 12%

average = tax / total taxable income

=10460/66000

=15.8%

effective tax rate= tax / (taxable income + tax exempt interest income)

=10460/(66000+8000)

=10460/74000

=14.1%

Answer C

12% marginal

15.8% average tax rate

14.1% effective tax rate